OpenStudy (anonymous):

Consider the following equation. 3x^2 - y^2 = 5 (a) Find y' by implicit differentiation. (b) Solve the equation explicitly for y and differentiate to get y' in terms of x.

OpenStudy (loser66):

Ah, I got it. derivative both sides, 6x - 2y y' =0 2yy' = 6 x y' = 6x/2y = 3x/y this is implicit derivative

OpenStudy (loser66):

this is explicit derivative: 3x^2 - y^2 =5 y^2 = 3x^2 -5 \[y = \pm\sqrt{3x^2 -5}\] take derivative both sides to get the answer.

OpenStudy (anonymous):

Thank you!